# Exploiting Agent Symmetries for Performance Analysis of Distributed Optimization Methods

### v1

@@ 5.1 The PEP formulation restricted to symmetric agent-class solutions
Based on the results from Proposition 6, the definitions of the symmetric PEP solutions $F^{s}_{\mathcal{T}}$ (105) and $G^{s}_{\mathcal{T}}$ (110), we now show when and how an agent-dependent PEP can be expressed in a compact form, using only the blocks $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,m$) as variables. This will enable efficient computing of $w(\mathcal{S}_{n})$ for all $n\geq 2$, even when all the agents are not equivalent. We remind that $w(\mathcal{S}_{n}=\{n,\mathcal{A},t,\mathcal{P},\mathcal{F}_{u},\mathcal{I},\mathcal{W}\})$ denotes the worst-case performance of the execution of $t$ iterations of distributed algorithm $\mathcal{A}$ with $n$ agents, with respect to the performance criterion $\mathcal{P}$, valid for any starting point satisfying the set of initial conditions $\mathcal{I}$, any local functions being each in a given set of function $f_{i}\in\mathcal{F}_{u_{i}}$ and any averaging matrix in the set of matrices $\mathcal{W}$.
@@
[…]
If the algorithm $\mathcal{A}\in\mathcal{A}_{D}$, the performance measure $\mathcal{P}$, the initial conditions $\mathcal{I}$, and the interpolation constraints for each $\mathcal{F}_{u}$ ($u=1,\dots,m$) are linearly (or LMI) Gram representable, then the computation of $w(\mathcal{S}_{n})$ can be formulated as a compact SDP PEP, with $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,m$) as variables and, hence, whose size only depends on the number of agent equivalence classes $m$ but not directly on the total number of agents $n$.
@@
The blocks $G_{B}^{u}$ contain scalar products between variables of different, but equivalent, agents. Therefore, $G_{B}^{u}$ does not exist for equivalence classes with only one agent $n_{u}=1$. The blocks $G_{E}^{uv}$ contain scalar products between variables of different agents that are not equivalent. Therefore, these blocks do not exist when $m=1$. The other blocks $f_{A}^{u}$ and $G_{A}^{u}$, related to only one agent, always exist.
@@
Let $G^{s}_{\mathcal{T}}\in\mathbb{R}^{np\times np}$ be a symmetrized agent-class Gram matrix, as defined in (110). The SDP constraint $G^{s}_{\mathcal{T}}\succeq 0$ can be expressed with $G_{A}^{u}\in\mathbb{R}^{p\times p}$, $G_{B}^{u}\in\mathbb{R}^{p\times p}$ and $G_{E}^{uv}\in\mathbb{R}^{p\times p}$, $u,v=1\dots,m$:
$G^{s}_{\mathcal{T}}\succeq 0\qquad\Leftrightarrow\qquad\begin{cases}G_{A}^{u}\succeq 0&\text{for all $u$ with $n_{u}=1$,}\\ G_{A}^{u}-G_{B}^{u}\succeq 0&\text{for all $u$ with $n_{u}\geq 2$,}\end{cases}\qquad\text{ and }\qquad H_{\mathcal{T}}\succeq 0,$ (111)
where $H_{\mathcal{T}}\in\mathbb{R}^{mp\times mp}$ is a symmetric matrix composed of $m^{2}$ blocks of dimension $p\times p$:
$H_{\mathcal{T}}^{uv}=\begin{cases}n_{u}G_{A}^{u}+n_{u}(n_{u}-1)G_{B}^{u},&\text{when $u=v$,}\\ n_{u}n_{v}G_{E}^{uv},&\text{when $u\neq v$}\end{cases}.\qquad u,v=1,\dots,m.$ (112)
@@
$G^{s}_{\mathcal{T}}\succeq 0\quad\Leftrightarrow z^{T}G^{s}_{\mathcal{T}}z\geq 0\quad\text{for all $z\in\mathbb{R}^{np}.$}$
The vector $z\in\mathbb{R}^{np}$ can be divided into $m$ sub-vectors of length $n_{u}p$ to match with the blocks of matrix $G^{s}_{\mathcal{T}}$ (see Proposition 6). We also decompose each sub-vector $z_{u}$ in two terms : the average of $z_{u}$ over the agents from the class $\mathcal{V}_{u}$, and the remaining part of $z^{\perp}_{u}$:
$z_{u}=\mathbf{1}_{n_{u}}\otimes\overline{z}_{u}+z^{\perp}_{u}\quad\text{ with }\quad\overline{z}_{u}=\frac{1}{n_{u}}(\mathbf{1}_{n_{u}}\otimes I_{p})^{T}z_{u},\qquad\overline{z}_{u}\in\mathbb{R}^{p},\leavevmode\nobreak\ z^{\perp}_{u}\in\mathbb{R}^{n_{u}p}\text{ and }z_{u}\in\mathbb{R}^{n_{u}p}.$
By definition, we have that $\frac{1}{n_{u}}(\mathbf{1}_{n_{u}}\otimes I_{p})^{T}z^{\perp}_{u}=0$ for all $j=1,\dots,m$. We now compute $z^{T}G^{s}_{\mathcal{T}}z$ by summing over each block:
$z^{T}G^{s}_{\mathcal{T}}z=\sum_{u=1}^{m}z_{u}^{T}G_{\mathcal{V}_{u}}^{s}z_{u}+\sum_{u=1}^{m}\sum_{v\neq u}z_{u}^{T}G_{\mathcal{V}_{u}\mathcal{V}_{v}}^{s}z_{v},$ (113)
@@
$z^{T}G^{s}_{\mathcal{T}}z=\overline{z}^{T}H_{\mathcal{T}}\overline{z}+\sum_{u=1}^{m}(z^{\perp}_{u})^{T}\quantity(I_{n_{u}}\otimes(G_{A}^{u}-G_{B}^{u}))z^{\perp}_{u},$
where $\overline{z}\in\mathbb{R}^{mp}$ is a vector stacking the average vectors $\overline{z}_{u}$ of each equivalence class ($u=1,\dots,m$), and $H_{\mathcal{T}}\in\mathbb{R}^{mp\times mp}$ is the matrix defined in (112). Therefore, the constraint $G^{s}_{\mathcal{T}}\succeq 0$ can be expressed as
[…]
Conditions $H_{\mathcal{T}}\succeq 0$ and $G_{A}^{u}-G_{B}^{u}\succeq 0$ (for all $u=1,\dots,m$) are sufficient to ensure that (119) is satisfied and hence that $G^{s}\succeq 0$. To show the necessity of these conditions, let us consider different cases for $\overline{z}$ and $z^{\perp}_{u}$:
  • When $z^{\perp}_{u}=0$ for all $u$, then (119) simply imposes that $\overline{z}^{T}H_{\mathcal{T}}\overline{z}\geq 0$ for all $\overline{z}\in\mathbb{R}^{mp}$, i.e. $H_{\mathcal{T}}\succeq 0$.
@@
    $\sum_{i=1}^{n_{v}}(z^{\perp}_{v}(i))^{T}(G_{A}^{v}-G_{B}^{v})z^{\perp}_{v}(i)\geq 0,\quad\text{for all $z^{\perp}_{v}\in\mathbb{R}^{p}$ such that $(\mathbf{1}_{n_{v}}\otimes I_{p})^{T}z^{\perp}_{v}=0$,}\vspace{-1mm}$
    where $z^{\perp}_{v}(i)\in\mathbb{R}^{p}$, for $i=1,\dots,n_{v}$, denotes the $i^{\mathrm{th}}$ part of vector $z^{\perp}_{v}\in\mathbb{R}^{n_{v}p}$. Since $n_{v}\geq 2$, we can choose $z^{\perp}_{v}(1)=-z^{\perp}_{v}(2)=s$, for an arbitrary $s\in\mathbb{R}^{p}$, and $z^{\perp}_{v}(j)=0$ for $j\geq 3$, so that the above equation becomes $s^{T}(G_{A}^{v}-G_{B}^{v})s\geq 0$ for all $s\in\mathbb{R}^{p}$, which is equivalent to $G_{A}^{v}-G_{B}^{v}\succeq 0$. This is valid for an arbitrary $v=1,\dots,m$.
@@
As detailed in Section 2.2, the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ has Gram-representable necessary interpolation constraints, see Theorem 4 applied to the consensus subspace $\mathcal{C}$. Since $\mathcal{A},\mathcal{P},\mathcal{I}$ and interpolation constraints for each $\mathcal{F}_{u}$ are also Gram-representable, Proposition 2 guarantees that the computation of $w(n,\mathcal{A},t,\mathcal{P},\mathcal{I},\mathcal{F}_{u},\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}})$ can be formulated as an agent-dependent SDP PEP problem (14). Moreover, by Proposition 6, we can restrict the PEP to symmetric agent-class solutions of the form of $F^{s}_{\mathcal{T}}$ (72) and $G^{s}_{\mathcal{T}}$ (76). The symmetric solutions $F^{s}_{\mathcal{T}}$ and $G^{s}_{\mathcal{T}}$ are composed of blocks $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,m$). Therefore, all the PEP elements can be expressed in terms of these smaller blocks. Lemma 9.1 shows how the SDP constraint $G^{s}_{\mathcal{T}}\succeq 0$, can be expressed in terms of $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$. The other elements of the PEP can directly be expressed in terms of $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,m$), using the definition of $F^{s}_{\mathcal{T}}$ (72) and $G^{s}_{\mathcal{T}}$ (81), as detailed in Appendix A. Since all the PEP elements can be expressed in terms of blocks $f_{A}^{u}\in\mathbb{R}^{q}$, $G_{A}^{u}\in\mathbb{R}^{p\times p}$, $G_{B}^{u}\in\mathbb{R}^{p\times p}$ (when $n_{u}\geq 2$) and $G_{E}^{uv}\in\mathbb{R}^{p\times p}$ ($u,v=1\dots,m$), the problem can be made smaller by considering them as the variables of the SDP PEP, instead of the full matrix $G^{s}_{\mathcal{T}}\in\mathbb{R}^{np\times np}$ and vector $F^{s}_{\mathcal{T}}\in\mathbb{R}^{nq}$. The size of the problem depends on the number $m$ of equivalence classes of agents, and not directly on the total number of agents. ∎
Theorem 9 shows that the worst-case performance of a distributed optimization method can be computed with an SDP PEP whose size only depends on the number $m$ of equivalence classes of agents, but not directly on the total number of agents $n$. Therefore, if the number of equivalence classes $m$ is independent of $n$, so is the size of the problem. However, the problem can depend on $n$ and $n_{u}$ ($u=1,\dots,m$) as parameters. Appendix A provides details to explicitly construct this compact PEP formulation.
@@
If the number $m$ of equivalence classes of agents is independent of $n$, so is the size of the compact SDP PEP, and we can solve it efficiently for any value of $n$. In particular, we can take the limit when $n$ tend to $\infty$ in the problem, to compute the worst-case performance of an algorithm on an infinitely large network of agents
@@
and to determine if this performance is bounded or not. In all the cases we treated (see Section 6), we observed that if the worst-case value for $n\to\infty$ is bounded, then it depends on the size proportions of each equivalence class $\rho_{1},\dots,\rho_{m}$:

### v2

@@ 5.1 The PEP formulation restricted to symmetric agent-class solutions
Based on the results from Proposition 6, the definitions of the symmetric PEP solutions $\mathbf{f}^{s}_{\mathcal{T}}$ (106) and $G^{s}_{\mathcal{T}}$ (107), we now show when and how an agent-dependent PEP can be expressed in a compact form, using only the blocks $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1,\dots,r$) as variables. This will enable efficient computing of $w(\mathcal{S}_{n})$ for all $n\geq 2$, even when all the agents are not equivalent. We remind that $w(\mathcal{S}_{n}=\{n,\mathcal{A},t,\mathcal{P},\mathcal{F}_{u},\mathcal{I},\mathcal{W}\})$ denotes the worst-case performance of the execution of $t$ iterations of distributed algorithm $\mathcal{A}$ with $n$ agents, with respect to the performance criterion $\mathcal{P}$, valid for any starting point satisfying the set of initial conditions $\mathcal{I}$, any local functions being each in a given set of function $f_{i}\in\mathcal{F}_{u_{i}}$ and any averaging matrix in the set of matrices $\mathcal{W}$.
@@
Let $\mathcal{V}$ be a set of $n\geq 2$ agents, $\mathcal{T}$ be a partition of $\mathcal{V}$ into $r\leq n$ equivalence classes of agents, as defined in Definition 8, and $\mathcal{W}=\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ with $\lambda^{-}\leq\lambda^{+}\in(-1,1)$.
If the algorithm $\mathcal{A}\in\mathcal{A}_{D}$, the performance measure $\mathcal{P}$, the initial conditions $\mathcal{I}$, and the interpolation constraints for each $\mathcal{F}_{u}$ ($u=1,\dots,r$) are linearly (or LMI) Gram representable, then the computation of $w(\mathcal{S}_{n})$ can be formulated as a compact SDP PEP, with $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,r$) as variables and, hence, whose size only depends on the number of agent equivalence classes $r$ but not directly on the total number of agents $n$.
@@
The blocks $G_{B}^{u}$ contain scalar products between variables of different, but equivalent, agents. Therefore, $G_{B}^{u}$ does not exist for equivalence classes with only one agent $n_{u}=1$. The blocks $G_{E}^{uv}$ contain scalar products between variables of different agents that are not equivalent. Therefore, these blocks do not exist when $r=1$. The other blocks $f_{A}^{u}$ and $G_{A}^{u}$, related to only one agent, always exist.
@@
Let $G^{s}_{\mathcal{T}}\in\mathbb{R}^{np\times np}$ be a symmetrized agent-class Gram matrix, as defined in (107). The SDP constraint $G^{s}_{\mathcal{T}}\succeq 0$ can be expressed with $G_{A}^{u}\in\mathbb{R}^{p\times p}$, $G_{B}^{u}\in\mathbb{R}^{p\times p}$ and $G_{E}^{uv}\in\mathbb{R}^{p\times p}$, $u,v=1\dots,r$ :
[…]
where $H_{\mathcal{T}}\in\mathbb{R}^{rp\times rp}$ is a symmetric matrix composed of $r^{2}$ blocks of dimension $p\times p$ :
$H_{\mathcal{T}}^{uv}=\begin{cases}n_{u}G_{A}^{u}+n_{u}(n_{u}-1)G_{B}^{u},&\text{when $u=v$,}\\ n_{u}n_{v}G_{E}^{uv},&\text{when $u\neq v$}\end{cases}.\qquad u,v=1,\dots,r.$ (109)
@@
$G^{s}_{\mathcal{T}}\succeq 0\quad\Leftrightarrow z^{T}G^{s}_{\mathcal{T}}z\geq 0\quad\text{for all vector $z\in\mathbb{R}^{np}.$}$
A vector $z\in\mathbb{R}^{np}$ can be divided into $r$ sub-vectors of length $n_{u}p$ to match with the blocks of matrix $G^{s}_{\mathcal{T}}$ (see Proposition 6). We also decompose each sub-vector $z_{u}$ in two terms : the average of $z_{u}$ over the agents from the class $\mathcal{V}_{u}$, and the remaining part of $z^{\perp}_{u}$ :
$z_{u}=\mathbf{1}_{n_{u}}\otimes\overline{z}_{u}+z^{\perp}_{u}\quad\text{ with }\quad\overline{z}_{u}=\frac{1}{n_{u}}(\mathbf{1}_{n_{u}}\otimes I_{p})^{T}z_{u},\qquad\overline{z}_{u}\in\mathbb{R}^{p},\penalty 10000\ z^{\perp}_{u}\in\mathbb{R}^{n_{u}p}\text{ and }z_{u}\in\mathbb{R}^{n_{u}p}.$
By definition, we have that $\frac{1}{n_{u}}(\mathbf{1}_{n_{u}}\otimes I_{p})^{T}z^{\perp}_{u}=0$ for all $j=1,\dots,r$. We now compute $z^{T}G^{s}_{\mathcal{T}}z$ (for any vector $z\in\mathbb{R}^{np}$) by summing over each block :
$z^{T}G^{s}_{\mathcal{T}}z=\sum_{u=1}^{r}z_{u}^{T}G_{\mathcal{V}_{u}}^{s}z_{u}+\sum_{u=1}^{r}\sum_{v\neq u}z_{u}^{T}G_{\mathcal{V}_{u}\mathcal{V}_{v}}^{s}z_{v},$ (110)
@@
$z^{T}G^{s}_{\mathcal{T}}z=\overline{z}^{T}H_{\mathcal{T}}\overline{z}+\sum_{u=1}^{r}(z^{\perp}_{u})^{T}\quantity(I_{n_{u}}\otimes(G_{A}^{u}-G_{B}^{u}))z^{\perp}_{u},$
where $\overline{z}\in\mathbb{R}^{rp}$ is a vector stacking the average vectors $\overline{z}_{u}$ of each equivalence class ($u=1,\dots,r$), and $H_{\mathcal{T}}\in\mathbb{R}^{rp\times rp}$ is the matrix defined in (109). Therefore, the constraint $G^{s}_{\mathcal{T}}\succeq 0$ can be expressed as
$\overline{z}^{T}H_{\mathcal{T}}\overline{z}+\sum_{u=1}^{r}(z^{\perp}_{u})^{T}\quantity(I_{n_{u}}\otimes(G_{A}^{u}-G_{B}^{u}))z^{\perp}_{u}\geq 0,\quad\text{\small{for all $\overline{z}\in\mathbb{R}^{rp}$ and all $z^{\perp}_{u}\in\mathbb{R}^{n_{u}p}$ s.t. $(\mathbf{1}_{n_{u}}\otimes I_{p})^{T}z^{\perp}_{u}=0$}}$ (116)
Conditions $H_{\mathcal{T}}\succeq 0$ and $G_{A}^{u}-G_{B}^{u}\succeq 0$ (for all $u=1,\dots,r$) are sufficient to ensure that (116) is satisfied and hence that $G^{s}\succeq 0$. To show the necessity of these conditions, let us consider different cases for $\overline{z}$ and $z^{\perp}_{u}$ :
  • When $z^{\perp}_{u}=0$ for all $u$, then (116) simply imposes that $\overline{z}^{T}H_{\mathcal{T}}\overline{z}\geq 0$ for all $\overline{z}\in\mathbb{R}^{rp}$, i.e. $H_{\mathcal{T}}\succeq 0$.
@@
    $\sum_{i=1}^{n_{v}}(z^{\perp}_{v}(i))^{T}(G_{A}^{v}-G_{B}^{v})z^{\perp}_{v}(i)\geq 0,\quad\text{for all $z^{\perp}_{v}\in\mathbb{R}^{p}$ such that $(\mathbf{1}_{n_{v}}\otimes I_{p})^{T}z^{\perp}_{v}=0$,}\vskip-2.84526pt$
    where $z^{\perp}_{v}(i)\in\mathbb{R}^{p}$, for $i=1,\dots,n_{v}$, denotes the $i^{\mathrm{th}}$ part of vector $z^{\perp}_{v}\in\mathbb{R}^{n_{v}p}$. Since $n_{v}\geq 2$, we can choose $z^{\perp}_{v}(1)=-z^{\perp}_{v}(2)=s$, for an arbitrary $s\in\mathbb{R}^{p}$, and $z^{\perp}_{v}(j)=0$ for $j\geq 3$, so that the above equation becomes $s^{T}(G_{A}^{v}-G_{B}^{v})s\geq 0$ for all $s\in\mathbb{R}^{p}$, which is equivalent to $G_{A}^{v}-G_{B}^{v}\succeq 0$. This is valid for an arbitrary $v=1,\dots,r$.
@@
As detailed in Section 2.2, the set of averaging matrices $\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}}$ has Gram-representable necessary interpolation constraints, see Corollary 4.1. Since $\mathcal{A},\mathcal{P},\mathcal{I}$ and interpolation constraints for each $\mathcal{F}_{u}$ are also Gram-representable, Proposition 2 guarantees that the computation of $w(n,\mathcal{A},t,\mathcal{P},\mathcal{I},\mathcal{F}_{u},\mathcal{W}_{\text{\footnotesize$(\lambda^{-},\lambda^{+})$}})$ can be formulated as an agent-dependent SDP PEP problem (15). Moreover, by Proposition 6, we can restrict the PEP to symmetric agent-class solutions of the form of $\mathbf{f}^{s}_{\mathcal{T}}$ (74) and $G^{s}_{\mathcal{T}}$ (75). The symmetric solutions $\mathbf{f}^{s}_{\mathcal{T}}$ and $G^{s}_{\mathcal{T}}$ are composed of blocks $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,r$). Therefore, all the PEP elements can be expressed in terms of these smaller blocks. Lemma 9.1 shows how the SDP constraint $G^{s}_{\mathcal{T}}\succeq 0$, can be expressed in terms of $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$. The other elements of the PEP can directly be expressed in terms of $f_{A}^{u}$, $G_{A}^{u}$, $G_{B}^{u}$, and $G_{E}^{uv}$ ($u,v=1\dots,r$), using the definition of $\mathbf{f}^{s}_{\mathcal{T}}$ (74) and $G^{s}_{\mathcal{T}}$ (77), as detailed in Appendix A. Since all the PEP elements can be expressed in terms of blocks $f_{A}^{u}\in\mathbb{R}^{q}$, $G_{A}^{u}\in\mathbb{R}^{p\times p}$, $G_{B}^{u}\in\mathbb{R}^{p\times p}$ (when $n_{u}\geq 2$) and $G_{E}^{uv}\in\mathbb{R}^{p\times p}$ ($u,v=1\dots,r$), the problem can be made smaller by considering them as the variables of the SDP PEP, instead of the full matrix $G^{s}_{\mathcal{T}}\in\mathbb{R}^{np\times np}$ and vector $\mathbf{f}^{s}_{\mathcal{T}}\in\mathbb{R}^{nq}$. The size of the problem depends on the number $r$ of equivalence classes of agents, and not directly on the total number of agents. ∎
Theorem 9 shows that the worst-case performance of a distributed optimization method can be computed with an SDP PEP whose size only depends on the number $r$ of equivalence classes of agents, but not directly on the total number of agents $n$. Therefore, if the number of equivalence classes $r$ is independent of $n$, so is the size of the problem. However, the problem can depend on $n$ and $n_{u}$ ($u=1,\dots,r$) as parameters. Appendix A provides details to explicitly construct this compact PEP formulation.
@@
If the number $r$ of equivalence classes of agents is independent of $n$, so is the size of the compact SDP PEP, and we can solve it efficiently for any value of $n$. In particular, we can take the limit when $n$ tend to $\infty$ in the problem, to compute the worst-case performance of an algorithm on an infinitely large network of agents
@@
and to determine if this performance is bounded or not. In all the cases we treated (see Section 6), we observed that if the worst-case value for $n\to\infty$ is bounded, then it depends on the size proportions of each equivalence class $\rho_{1},\dots,\rho_{r}$ :